OpenStudy (anonymous):

Just another cute problem: Find the range of \(x\) for which \( \arccos \sqrt{x-1}- \arcsin \sqrt{2-x} =0 \) holds.

OpenStudy (anonymous):

x => 1 And x <= 2 x \(\in\) [1,2]

OpenStudy (blockcolder):

Take cos of both sides: \[\sqrt{x-1}=\cos({\arcsin{\sqrt{2-x}})}\]

OpenStudy (anonymous):

@blockcolder: That's not how I did it. @Ishaan94: How you did it?

OpenStudy (blockcolder):

cos(arcsin(x))=sqrt(1-x^2), so \[\sqrt{x-1}=1-(2-x)=1+x\] Then solve this for x.

OpenStudy (anonymous):

\[\sqrt{x-1}=1-(2-x) \neq1+x\]

OpenStudy (blockcolder):

Oops. Typo. =)) \[\sqrt{x-1}=x-1\] Only when x-1=0 or x-1=1 which means x=1 or 2.

OpenStudy (anonymous):

Check again ^^

OpenStudy (anonymous):

I got the domain for square-root part. And then checked if it agrees with the domain of inverse functions.

OpenStudy (anonymous):

Interesting approach Ishaan. +1.